OpenStudy (anonymous):

Question on Infinite Limit, involving a Radical. lim x→∞√(4x^2+3x) −2x

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} \sqrt{4x ^{2} + 3x} -2x\]

OpenStudy (anonymous):

\[\lim_{x\to\infty}\left(\sqrt{4x^2+3x}-2x\right)\] \[\sqrt{4x^2+3x}-2x\cdot\frac{\sqrt{4x^2+3x}+2x}{\sqrt{4x^2+3x}+2x}\\ \frac{4x^2+3x-4x^2}{\sqrt{4x^2+3x}+2x}\\ \frac{3x}{\sqrt{4x^2+3x}+2x}\\ \frac{3x}{\sqrt{x^2}\sqrt{4+\frac{3}{x}}+2x}\\ \frac{3x}{|x|\sqrt{4+\frac{3}{x}}+2x}\] Since \(x\to\infty\), you have \(|x|=x\): \[\frac{3x}{x\sqrt{4+\frac{3}{x}}+2x}\\ \frac{3}{\sqrt{4+\frac{3}{x}}+2}\] Taking the limit is easy from here.

OpenStudy (anonymous):

Thank you.